# Denoising Diffusion Probabilistic Models for Coastal Inundation Forecasting

### v1

@@ Abstract
Coastal flooding poses significant risks to communities, necessitating fast and accurate forecasting methods to mitigate potential damage. To approach this problem, we present Diff-Flood, a probabilistic spatiotemporal forecasting method designed based on denoising diffusion models. Diff-Flood predicts inundation level at a location by taking both spatial and temporal context into account. It utilizes inundation levels at neighboring locations and digital elevation data as spatial context. Inundation history from a context time window, together with additional co-variates are used as temporal context. Convolutional neural networks and cross-attention mechanism are then employed to capture the spatiotemporal dynamics in the data. We trained and tested Diff-Flood on coastal inundation data from the Eastern Shore of Virginia, a region highly impacted by coastal flooding. Our results show that, Diff-Flood outperforms existing forecasting methods in terms of prediction performance ($6\%\text{ to }64\%$ improvement in terms of two performance metrics) and scalability.
1 Introduction
@@
The projection in sea level rise have highlighted the increasing vulnerability of many coastal regions to inundation (Wuebbles et al., 2017). Among these regions, the East Coast in US is a region of particular policy concern due to its geography and low-lying topography (Ezer and Atkinson, 2014). This region, in recent periods, has been experiencing frequent inundation events  (Ezer, 2020), causing disruptions in daily life (e.g. traffic, economic) (Jacobs et al., 2018). In the future, this situation will only keep declining due to the combined effect of climate change, sea level rise, and urbanization (Swain et al., 2020). Therefore, forecasting coastal inundation is of utmost importance from a policy standpoint for undertaking risk prevention measurement or safe evacuation of the residents in danger  (Yang et al., 2019).
Figure 1: Flooding on Quinby Bridge Road, Accomack County, VA on April 3, 2024 (Data source: Tidewatch Maps (Tidewatch, 2025).)
The problem of coastal inundation forecasting came to our attention during conversations with staff at the City of Virginia Beach. The region faces challenges with coastal flooding which causes slow-down and/or road closures; disrupting the day-to-day activities of the inhabitants. A fast and high-resolution flood forecasting model would enable city officials to identify localized inundation risks, resulting in more effective preparedness and response. In this paper, we approach the problem of designing such a model. It is a joint-work of a multidisciplinary team, including computer and data scientists, academic experts with expertise in hydro-dynamic models, members of established coastal resources management centers, and officials at the City of Virginia Beach.
@@
The traditional method for predicting inundation involves hydro-dynamical, physics-based models (PBM). Commercially available (Syme, 2001; Deltares, 2014) (e.g., SOBEK, TUFLOW) or open-sourced (Zhang et al., 2016) (e.g, SCHISM) hydro-dynamical models can simulate inundation accurately at a high resolution. However, the computation time associated with these PBMs makes them challenging for real-time forecasting (Guo and others, 2021). Although researchers have investigated Deep-Learning (DL) based time-series prediction methods as surrogates to these PBMs (Roy et al., 2023), these models are not suitable for accommodating spatial contexts. Existing spatiotemporal forecasting methods, on the other hand, struggle to scale in high-resolution forecasting tasks.
@@
Motivated by the above observations, we explore the Diffusion Model (Sohl-Dickstein et al., 2015) in the context of high-resolution coastal inundation forecasting task. The primary reason for investigating diffusion models is its ability to utilize conditioning contexts of various dimensions (Rombach et al., 2022; Zhang and others, 2023) and its potential to model physical properties (Ahmad and others, 2024; Amram and Pedro, 2023). Originally developed for image generation, diffusion models have been adapted for time-series forecasting, achieving state-of-the-art performance, as demonstrated by Rasul et al. (2021). However, their approach does not incorporate spatial context. More recently,  (Wen et al., 2023) proposed a probabilistic spatiotemporal graph forecasting method based on denoising diffusion models. While promising, such methods face scalability challenges when applied to large-scale/high-resolution inundation forecasting tasks. Using the coastal inundation of the Eastern Shore as case study, we address these challenges by making the following contributions in this paper.
  • We formulate the problem of coastal inundation forecasting under the spatiotemporal forecasting framework and introduce Diff-Flood as a novel method for approaching the problem. It can utilize both spatial and temporal context to generate spatiotemporal forecasts at fine spatial resolution.
  • Extensive evaluation of our method against several baseline methods on inundation data of the Virginia Eastern Shore underscores the superiority of our proposed method and its aspect of being used as a decision-making tool for policymakers. Specifically, we observed $6\%$ to $64\%$ improvement in prediction performance, in terms of two performance metrics (Normalized Average Continuous Ranked Probability Score, and Normalized Root Mean Square Error). An ablation study of our model is performed to highlight the importance of its components.
  • Furthermore, we demonstrate the scalability of our model compared to other baselines in large settings, underscoring its utility in real-time forecasting in large-scale scenarios.
It is worth highlighting that our model can be used to answer inundation scenario-based queries relevant to policymakers. For example, queries of the form: “What is the probability that a particular region will get inundated by 1 feet within the next 12 hours?”, can be answered by using our model through probabilistic forecasting, see section 7.
2 Related Work
@@
Hydrology Models: Traditionally, various Physics Based hydrodynamic Models (PBM) have been employed to simulate inundation. SOBEK (Deltares, 2014) is a 2D model that solves a Saint-Venant flow equation and shallow water equation to determine the water level of rectangular grids. TELEMAC-2D (Vu et al., 2015) uses a similar method but is more suitable for coastal areas. Polymorphic models like SCHISM (Zhang et al., 2016) can capture water flow at very high resolution. However, the computation time associated with solving these complex equations at high spatial resolution makes these PBMs infeasible for real-time forecasting.
Deep Learning: DL methods are a popular choice in surrogating hydrodynamic models. Regression-based methods (Zahura and Goodall, 2022), recurrent neural networks (Roy et al., 2023) and hybrid ML-based methods (Zanchetta and Coulibaly, 2022) are some of the methods that have been attempted as possible surrogates. However, these models cannot inherently take spatial context into account. The DCRNN method proposed by (Li et al., 2017) for traffic forecasting is the state-of-the-art DL method for spatiotemporal point forecasting, where bidirectional random walk is used to capture spatial dependency and Seq2Seq architecture is used to capture temporal dependency.
@@
Diffusion Model The field of generative modeling has been dominated by the diffusion model for quite some time. Although it had primarily been studied for image synthesis (Ho et al., 2020; Dhariwal and Nichol, 2021; Austin et al., 2021), its fascinating ability to use flexible conditioning to guide the generation process has been utilized in other domains like video generation (Yang et al., 2023b), prompt-based image generation (Ramesh et al., 2022), text-to-speech generation (Yang et al., 2023a). TimeGrad (Rasul et al., 2021) was the pioneering method that approached the time-series forecasting problem by utilizing a diffusion model. Other models like CSDI (Tashiro et al., 2021), DPSD-CSPD (Biloš et al., 2023), TDSTF (Chang et al., 2024) were later introduced. These models try to predict the noise added during the forward process of the diffusion model ($\epsilon$-parameterization). An alternative parameterization is to predict the ground-truth data ($x$-parameterization). Different from models mentioned above, $D^{3}$ (Li et al., 2022) uses coupled diffusion process and bidirectional variational autoencoder to better forecast in longer horizon and improve interpretability. The State-of-the-Art Diffusion-based model DiffSTG (Wen et al., 2023) devised a modified architecture for the reverse process of diffusion model combining UNet and GNN to capture temporal and spatial dependencies, respectively. However, both these methods scale poorly with the growth in number of variables.
3 Problem Formulation
@@
Formally, we define the inundation forecasting problem as follows.
Problem 1.
[…]
Problem 2.
@@
Multi-patch Probabilistic Inundation Forecasting Let $\mathcal{P}=\{P_{1},P_{2},...,P_{K}\}$ denote a set of two-dimensional patches of size ${D\times D}$. Let ${}_{k}\mathbf{x}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the inundation matrix of patch $P_{k}$ at timestep $t\in\{1,\ldots,T\}$. Let ${}_{k}\mathbf{s}\in\mathbb{R}^{D\times D}$ and ${}_{k}\mathbf{z}_{t}\in\mathbb{R}^{l},t\in\{1,\ldots,T\}$ denote the elevation matrix and time series of co-variates associated with patch $P_{k}$, respectively. Let $t_{0}:t_{c+l}$ denote the increasing sequence of timesteps ${t_{0},t_{1},...,t_{c+l}}$, where $c,l>0$. We aim to learn the conditional distribution function $q({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{x}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t},{}_{k}\mathbf{s})$.
Note that, both problems are spatio-temporal forecasting problems, since we aim to forecast the inundation level on all cells in a patch together. They are also probabilistic forecasting problems, as we want to learn the distribution of inundation values instead of making a single prediction. Finally, Problem 2 represents a generalized version of Problem 1. Here, we aim to learn a single model that can be utilized to forecast inundation levels on multiple patches/areas.
@@
The following section presents the basics of a denoising diffusion model used by our method Diff-Flood for learning our target distributions. Subsequently, we present details on how we have designed Diff-Flood to capture spatial and temporal dynamics of our inundation values.
4 Denoising Diffusion Model
@@
In this section, we provide a brief overview of denoising diffusion models. A more comprehensive view is provided in the supplementary material.
[…]
@@
[…]
Here, we get (2) by applying Jensen’s inequality. We indirectly minimize the negative log-likelihood of the data, i.e., $-\log p_{\theta}(\mathbf{x}^{0})$, by minimizing the right hand side of (3).
Ho et al. (2020) showed that the forward process has the following property: we can sample $\mathbf{x}^{n}$ using $\mathbf{x}^{0}$ at any noise step $n$ in closed form . Let, $\alpha_{n}=1-\beta_{n}$, and $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{i}$. Then, we have:
@@
Based on this property, Ho et al. (2020) showed that one possible parameterization of (1) is as follows:
@@
Here, $\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)$ is a neural network that predicts $\mathbf{x}^{0}$ from noisy vector $\mathbf{x}^{n}$ and noise step $n$. Ho et al. (2020) then showed that, the right hand side of (3) can be minimized by minimizing:
$\mathbb{E}_{n,q(\mathbf{x}^{n}|\mathbf{x}^{0})}\ \frac{1}{2\tilde{\beta}_{n}}\frac{\bar{\alpha}_{n-1}\ \beta_{n}^{2}}{(1-\bar{\alpha}_{n})^{2}}||\mathbf{x}^{0}-\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)||_{2}^{2}$ (7)
@@
$\mathbf{x}^{n-1}=\mu_{\theta}(\mathbf{x}^{n},n)+\sqrt{\Sigma_{\theta}}\ \mathbf{v}$ (8)
Here, $\mathbf{v}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$ for $n\in[2,N]$, and $\mathbf{v}={\mathbf{0}}$ when $n=1$. $\mu_{\theta}(\mathbf{x}^{n},n)$ and $\Sigma_{\theta}$ are computed using (5) and (6), respectively.
5 Diff-Flood Method
Diff-Flood is designed for multi-patch probabilistic inundation forecasting, where the inundation values are correlated both spatially and temporally. Following our notations in Problem 1 and 2, let ${}_{k}\mathbf{x}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the inundation matrix of the patch $P_{k}$ at time step $t$. Towards our goal of learning the conditional distribution $q_{\chi}({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{x}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t_{0}:t_{c-1}},{}_{k}\mathbf{s})$, where ${}_{k}\mathbf{s}\in\mathbb{R}^{D\times D}$ is the elevation matrix of patch $P_{k}$, and ${}_{k}\mathbf{z}_{t}$ denotes a time-series of co-variates associated with $P_{k}$, we assume:
$q_{\chi}({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{x}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t_{0}:t_{c-1}},{}_{k}\mathbf{s})=$
$\quad\quad\quad\textstyle\prod_{t=t_{c}}^{t_{c+l}}q_{\chi}({}_{k}\mathbf{x}_{t}^{0}|{}_{k}\mathbf{x}_{t_{0}:t-1}^{0},{}_{k}\mathbf{z}_{t_{0}:t-1},{}_{k}\mathbf{s})$ (9)
In Diff-Flood, we use a denoising diffusion model (as described in Section B) to learn the conditional distributions on the right hand side of (9). We refer to ${}_{k}\mathbf{x}_{t_{0}:t-1}^{0}$, ${}_{k}\mathbf{z}_{t_{0}:t-1}$, and ${}_{k}\mathbf{s}$ as context data.
Figure 2: Architecture of proposed Diff-Flood Model
To capture spatial correlation, we first augment the inundation data in the context (i.e. ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) by concatenating the elevation matrix ${}_{k}\mathbf{s}$ to it as an additional channel (Equation 10). The goal is to capture the correlation of elevation and inundation values. We then use convolution blocks extract spatial features ${}_{k}\mathbf{f}_{t_{0}:t-1}$ (Equation 11).
${}_{k}\mathbf{y}^{0}_{t_{0}:t-1}={}_{k}\mathbf{x}^{0}_{t_{0}:t-1}\oplus{}_{k}\mathbf{s}$ (10)
${}_{k}\mathbf{f}_{t_{0}:t-1}=\textsc{conv}{}_{\theta_{1}}({}_{k}\mathbf{y}^{0}_{t_{0}:t-1})$ (11)
In addition, since coastal inundation is expected to be correlated with tide cycles, we use the hour of day, and day of month as temporal features $\mathbf{z}_{t}$. We use sinusoidal encoding to capture the cyclic nature of these temporal features.
${}_{k}\mathbf{g}^{0}_{t_{0}:t-1}=\textsc{sinusoid}{}({}_{k}\mathbf{z}_{t_{0}:t-1})$ (12)
@@ 5 Diff-Flood Method
${}_{k}\mathbf{h}_{t_{0}:t-1}=\textsc{linear}{}_{\theta_{2}}({}_{k}\mathbf{f}_{t_{0}:t-1}\oplus{}_{k}\mathbf{g}_{t_{0}:t-1})$ (13)
An overview of the architecture of Diff-Flood is provided in Figure 2. We approximate (9) by the model in (14). Here, $\theta$ comprises of the trainable parameters of the convolution blocks ($\theta_{1}$), fully connected layer ($\theta_{2}$), and the denoising diffusion model ($\theta_{3}$).

### v2

# Towards High Resolution Probabilistic Coastal Inundation Forecasting from Sparse Observations

@@ Abstract
Coastal flooding poses increasing threats to communities worldwide, necessitating accurate and hyper-local inundation forecasting for effective emergency response. However, real-world deployment of forecasting systems is often constrained by sparse sensor networks, where only a limited subset of locations may have sensors due to budget constraints. To approach this challenge, we present Diff-Sparse, a masked conditional diffusion model designed for probabilistic coastal inundation forecasting from sparse sensor observations. Diff-Sparse primarily utilizes the inundation history of a location and its neighboring locations from a context time window as spatiotemporal context. The fundamental challenge of spatiotemporal prediction based on sparse observations in the context window is addressed by introducing a novel masking strategy during training. Digital elevation data and temporal co-variates are utilized as additional spatial and temporal contexts, respectively. A convolutional neural network and a conditional UNet architecture with cross-attention mechanism are employed to capture the spatiotemporal dynamics in the data. We trained and tested Diff-Sparse on coastal inundation data from the Eastern Shore of Virginia and systematically assessed the performance of Diff-Sparse across different sparsity levels ($0\%,50\%,95\%$ missing observations). Our experiment results show that Diff-Sparse achieves upto $62\%$ improvement in terms of two forecasting performance metrics compared to existing methods, at $95\%$ sparsity level. Moreover, our ablation studies reveal that digital elevation data becomes more useful at high sparsity levels compared to temporal co-variates.
Codehttps://github.com/KAI10/Diff-Sparse
Introduction
@@
The projection in sea level rise have highlighted the increasing vulnerability of many coastal regions to inundation (wuebbles2017climate). Among these regions, the East Coast in US is a region of particular policy concern due to its geography and low-lying topography (ezer2014accelerated). In recent periods, this region has been experiencing frequent inundation events  (ezer2020analysis), causing disruptions in daily life (e.g., traffic, economic) (jacobs2018recent). In future, this situation will only continue to decline due to the combined effect of climate change, sea level rise, and urbanization (swain2020increased). Therefore, forecasting coastal inundation is of utmost importance from a policy standpoint for undertaking risk prevention measurement and safe evacuation of residents during emergencies  (yang2019incorporating).
The problem of forecasting coastal inundation from sparse observation was brought to our attention through conversations with staff at the City of Virginia Beach. The region faces challenges with coastal flooding which causes slow-down and/or road closures; disrupting the day-to-day activities of the inhabitants. A fast and high-resolution flood forecasting model would enable city officials to identify localized flood risks, resulting in more effective preparedness and response. In order to effectively identify inundated locations, policymakers deploy sensors at various locations (tao2024review). These sensors are responsible for providing periodical inundation reports and other variables of interest in the surrounding area. However, there are several challenges associated with these sensor placements. First, due to budget and resource constraints, the number of sensors that can be deployed is limited (saad2024scalable; karagulian2019review). Second, the readings from the sensor are often associated with noise due to mechanical, environmental or calibration issues (barrenetxea2008assessing; mydlarz2024floodnet). Third, due to the difficulty associated with finding an optimal location assignment of spatial sensors (krause2008near) or changes carried out by the stakeholders, the locations of these sensors may change from time to time (garg2012learning). As a result, a forecasting method that relies on sensor data but does not account for the sparsity and uncertain nature of sensor placements will not be reliable (tao2024review).
@@
The traditional method for predicting inundation involves hydro-dynamical, physics-based models (PBM). Commercially available (syme2001tuflow; deltares2014sobek) (e.g., SOBEK, TUFLOW) or open-sourced (zhang2016seamless) (e.g, SCHISM) hydro-dynamical models can simulate inundation accurately at a high resolution. However, the computation time associated with these PBMs makes them challenging for real-time forecasting (guo2021urban). Researchers have investigated the use of Deep-Learning (DL) based prediction methods as surrogates to these PBMs (roy2023application). However, these models often struggle to: (i) scale in large-scale high-resolution forecasting tasks under sparse availability of data, and (ii) effectively use additional spatiotemporal contexts. Therefore, a surrogate model to PBMs that can make accurate real-time predictions based on sparse sensor observations, will have high relevance to policymakers.
@@
Motivated by the above observations, we explore Diffusion Models (sohl2015deep) for high-resolution coastal inundation forecasting task. The reason is two-fold: their potential to model physical properties (ahmad2024calobench; amram2023denoising), and their ability to utilize conditioning contexts of various dimensions (Rombach_2022_CVPR; zhang2023adding) that can improve prediction quality in the presence of sparse observations. Although diffusion models were originally developed for image generation, they have been adapted for time-series and spatiotemporal forecasting (rasul2021autoregressive; wen2023diffstg). However, the existing methods struggle in terms of computational scalability and performance, under sparse availability of data. Using the coastal inundation of the Virginia Eastern Shore as case study, we address these challenges by making the following contributions:
  • We formulate the problem of probabilistic inundation forecasting from sparse observations as a conditional distribution function learning problem and present Diff-Sparse, a scalable model to learn this function. Diff-Sparse is trained using data from physics-based hydro-dynamic model; however, during inference it makes probabilistic forecasts based on sparse sensor observations and additional spatiotemporal contexts (e.g., elevation, temporal co-variates).
  • We introduce a novel masking strategy to train Diff-Sparse which enables it to make forecasts from sparse observations. Our robust training strategy enables Diff-Sparse to make accurate spatiotemporal forecasts from different spatial placement configurations of the sensors, without requiring training the model from scratch.
  • Extensive evaluation of our method against several baseline methods under sparse observation underscores the superiority of our method at large-scale high resolution inundation forecasting tasks. Specifically, we observe upto $62\%$ improvement in prediction performance over existing methods in terms of two predictive performance metrics.
  • We perform extensive ablation studies to highlight the utility of different context data (e.g., sensor data, elevation, temporal co-variates) at varying sparsity levels.
Related Work
@@
Physcis-based Models: Traditionally, various Physics-Based hydrodynamic Models (PBMs) have been employed to simulate inundation. SOBEK (deltares2014sobek) is a two-dimensional model that solves the Saint-Venant flow equations and shallow water equations to determine water levels across rectangular grids. Similarly, TELEMAC-2D (vu2015two) employs comparable method specifically tailored for coastal applications. Polymorphic models such as SCHISM (zhang2016seamless) captures water flow at very high resolutions. However, the substantial computational time associated with solving these complex equations at high spatial resolution renders these PBMs impractical for real-time forecasting applications.
Deep Learning Researchers have explored Deep Learning (DL) methods for developing surrogates to PBMs. Regression-based methods (zahura2022predicting), recurrent neural networks (roy2023application), and hybrid machine learning-based methods (zanchetta2022hybrid) have been emerged as popular choices, among others. However, these models inherently struggle to incorporate spatiotemporal contexts effectively. Originally developed for traffic forecasting, the Diffusion Convolutional Recurrent Neural Network (DCRNN) method proposed by Li et al. li2017diffusion captures some aspect of spatiotemporal contexts by utilizing bidirectional random walks and Sequence-to-Sequence architecture. More recently, BayesNF (saad2024scalable) has emerged as the state-of-the-art DL spatiotemporal forecasting model, constructed using a Bayesian Neural Network framework. By assigning prior distributions to model parameters and adjusting the posterior based on observations, BayesNF effectively learns to map multivariate spatiotemporal points to continuous real-valued fields.
@@
Diffusion Model The field of generative modeling has been dominated by the diffusion model for quite some time. Although it had primarily been studied for image synthesis (ho2020denoising; dhariwal2021diffusion; austin2021structured), its fascinating ability to use flexible conditioning to guide the generation process has been utilized in other domains like video generation (yang2023diffusion), prompt-based image generation (ramesh2022hierarchical), text-to-speech generation (yang2023diffsound). TimeGrad (rasul2021autoregressive) was the pioneering method that approached the time-series forecasting problem by utilizing a diffusion model. Other models like CSDI (tashiro2021csdi), DPSD-CSPD (bilovs2023modeling), TDSTF (Chang2024) were later introduced. These models try to predict the noise added during the forward process of the diffusion model ($\epsilon$-parameterization). An alternative parameterization is to predict the ground-truth data ($x$-parameterization). Different from models mentioned above, $D^{3}$VAE (li2022generative) uses coupled diffusion process and bidirectional variational autoencoder to better forecast in longer horizon and improve interpretability. The State-of-the-Art Diffusion-based model DiffSTG (wen2023diffstg) devised a modified architecture for the reverse process of diffusion model combining UNet and GNN to capture temporal and spatial dependencies, respectively. However, both methods scale poorly with the growth of variables.
Problem Formulation
@@
Formally, we define the inundation forecasting from sparse observations problem as follows.
Problem 1
@@
Multi-patch Probabilistic Inundation Forecasting. Let $\mathcal{P}=\{P_{1},P_{2},...,P_{K}\}$ denote a set of two-dimensional rectangular grids each of size $D\times D$. We refer to $P_{k}$ as the $k^{th}$ ‘patch’. Let $\mathcal{M}=\{\mathbf{M}_{1},\mathbf{M}_{2},...,\mathbf{M}_{K}\}$ denote a set of binary matrices, henceforth referred to as ‘sensor masks’. If $\mathbf{M}_{k}(i,j)=1$ then the cell $(i,j)$ of patch $P_{k}$ has a water-level sensor; $\mathbf{M}_{k}(i,j)=0$ means cell $(i,j)$ of patch $P_{k}$ does not have a water-level sensor. Let $_{k}\mathbf{y}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the inundation matrix of patch $P_{k}$ at timestep $t$. If $\mathbf{M}_{k}(i,j)=1$ then $_{k}\mathbf{y}_{ij,t}^{0}\in\mathbb{R}$ contains the inundation level on cell $(i,j)$ of patch $P_{k}$ at time step $t$; $i,j\in\left\{1,...,D\right\},t\in\{1,...,T\}$. However, for cells $(i,j)$ where $\mathbf{M}_{k}(i,j)=0$, $_{k}\mathbf{y}_{ij,t}^{0},t\in\{1,...,T\}$ has missing value. Let $_{k}\mathbf{x}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the complete inundation matrix of patch $P_{k}$ at timestep $t$, where there are no missing values. Let $\mathbf{s}_{k}\in\mathbb{R}^{D\times D}$ and $_{k}\mathbf{z}_{t}\in\mathbb{R}^{l},t\in\{1,\ldots,T\}$ denote the elevation matrix and time series of co-variates associated with patch $P_{k}$, respectively. Let $t_{0}:t_{c+l}$ denote the increasing sequence of timesteps $t_{0},t_{1},...,t_{c+l}$}, where $c,l>0$. We aim to learn the conditional distribution function $q({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{y}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t},\mathbf{s}_{k},\mathbf{M}_{k})$.
Here, we aim to forecast the inundation on all cells in a patch (i.e., $\mathbf{x}$ ) even though in the context inundation history (i.e., $\mathbf{y}$) we only know the inundation levels on cells where sensors are placed. Problem 1 is a probabilistic forecasting problem, as we want to learn the distribution of inundation values instead of making a single prediction.
@@
The following section presents the basics of a denoising diffusion model used by our method Diff-Sparse for learning our target distributions. Subsequently, we present details on how we have designed Diff-Sparse to capture spatial and temporal dynamics of inundation values.
Denoising Diffusion Model
@@
In this section, we provide a brief overview of denoising diffusion models. A more comprehensive view is provided in the supplementary materials.
Let $\mathbf{x}^{0}\sim q_{\chi}(\mathbf{x}^{0})$ denote the multivariate training vector from some input space $\chi=\mathbb{R}^{D}$, where $q_{\chi}(\mathbf{x}^{0})$ denotes the true distribution. In diffusion probabilistic models, we aim to approximate $q_{\chi}(\mathbf{x}^{0})$ by a probability density function $p_{\theta}(\mathbf{x}^{0})$ where $\theta$ denotes the set of trainable parameters. Diffusion probabilistic models (sohl2015deep; luo2022understanding) are a special class of Hierarchical Variational Autoencoders (kingma2016improved; sonderby2016ladder) with the form $p_{\theta}(\mathbf{x}^{0})=\int p_{\theta}(\mathbf{x}^{0:N})d\mathbf{x}^{1:N}$; here $\mathbf{x}^{1},\mathbf{x}^{2},...,\mathbf{x}^{N}$ are latent variables. Three key properties of diffusion models are: $(i)$ all of the latent variables are assumed to have the dimension $D$ as $\mathbf{x}^{0}$, $(ii)$ the approximate posterior ,
@@
is fixed to a Markov chain (often referred to as the forward process), and $(iii)$ the structure of the latent encoder at each hierarchical step is pre-defined as a linear Gaussian model:
@@
$-\log p_{\theta}(\mathbf{x}^{0})\leq\mathbb{E}_{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}\log\frac{q(\mathbf{x}^{1:N}|\mathbf{x}^{0})}{p_{\theta}(\mathbf{x}^{0:N})}$ (2)
We indirectly minimize the negative log-likelihood of the data by minimizing the right hand side of (2).
(ho2020denoising) showed that the forward process has the following property: we can sample $\mathbf{x}^{n}$ using $\mathbf{x}^{0}$ at any noise step $n$ in closed form . Let, $\alpha_{n}=1-\beta_{n}$, and $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{i}$. Then, we have:
@@
Based on this property, (ho2020denoising) showed that one possible parameterization of (1) is as follows:
@@
Here, $\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)$ is a neural network that predicts $\mathbf{x}^{0}$ from noisy vector $\mathbf{x}^{n}$ and noise step $n$. (ho2020denoising) then showed that, the right hand side of (2) can be minimized by minimizing:
[…]
@@
$\mathbf{x}^{n-1}=\mu_{\theta}(\mathbf{x}^{n},n)+\sqrt{\Sigma_{\theta}}\ \mathbf{u}$ (7)
Here, $\mathbf{u}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$ for $n\in[2,N]$, and $\mathbf{u}={\mathbf{0}}$ when $n=1$. $\mu_{\theta}(\mathbf{x}^{n},n)$, $\Sigma_{\theta}$ are computed using (4) and (5), respectively.
Diff-Sparse Method
Diff-Sparse is designed for probabilistic inundation forecasting from sparse observations. Following our notations in Problem 1, let $_{k}\mathbf{y}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the ‘sparse’ inundation matrix of the patch $P_{k}$ at time step $t$, with sensor mask $\mathbf{M}_{k}$. Let $_{k}\mathbf{x}_{t}^{0}\in\mathbb{R}^{D\times D}$ denote the complete inundation matrix of patch $P_{k}$ at timestep $t$. Towards our goal of learning the conditional distribution $q_{\chi}({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{y}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t_{0}:t_{c-1}},\mathbf{s}_{k},\mathbf{M}_{k})$, where $\mathbf{s}_{k}\in\mathbb{R}^{D\times D}$ is the elevation matrix of patch $P_{k}$, and $_{k}\mathbf{z}_{t}$ denotes a time-series of co-variates associated with $P_{k}$, we assume:
$q_{\chi}({}_{k}\mathbf{x}_{t_{c}:t_{c+l}}^{0}|{}_{k}\mathbf{y}_{t_{0}:t_{c-1}}^{0},{}_{k}\mathbf{z}_{t_{0}:t_{c-1}},\mathbf{s}_{k},\mathbf{M}_{k})=$
$\quad\quad\quad\textstyle\prod_{t=t_{c}}^{t_{c+l}}q_{\chi}({}_{k}\mathbf{x}_{t}^{0}|{}_{k}\mathbf{y}_{t_{0}:t-1}^{0},{}_{k}\mathbf{z}_{t_{0}:t-1},\mathbf{s}_{k},\mathbf{M}_{k})$ (8)
In Diff-Sparse, we use a denoising diffusion model to learn the conditional distributions on the right hand side of (8). We refer to $_{k}\mathbf{y}_{t_{0}:t-1}^{0}$, $_{k}\mathbf{z}_{t_{0}:t-1}$, $\mathbf{s}_{k}$, and $M_{k}$ together as context data.
Figure 1: Diff-Sparse architecture. It has two main components. The first component involves context learning, where the sparse context inundation data ($\mathbf{y}_{t_{0}},\mathbf{y}_{t_{1}},...,\mathbf{y}_{t_{c-1}}$), sensor mask ($\mathbf{M}$), elevation data ($\mathbf{s}$) and temporal covariates ($\mathbf{z}_{t_{0}:t_{c-1}}$) are used to generate context embedding $\mathbf{h}_{t_{0}:t_{c-1}}$. The second component involves a diffusion model where the corresponding forecast is sampled by conditioning on the context embedding. $q(\mathbf{x}_{t_{c}}^{n}|\mathbf{x}_{t_{c}}^{n-1})$ denotes the forward diffusion process; $p_{\theta}(\mathbf{x}_{t_{c}}^{n-1}|\mathbf{x}_{t_{c}}^{n},n,\mathbf{h}_{t_{0}:t_{c-1}})$ denotes the reverse diffusion process, parameterized by $\theta$.
To capture spatial correlation, we augment the sparse inundation data in the context (i.e., $_{k}\mathbf{y}^{0}_{t_{0}:t-1}$) by concatenating the elevation matrix $\mathbf{s}_{k}$ and the sensor mask $\mathbf{M}_{k}$ to it as additional channels (9). The goal is to capture the correlation of elevation and inundation values. We then use convolution blocks to extract spatial features $_{k}\mathbf{f}_{t_{0}:t-1}$ (10).
${}_{k}\mathbf{v}_{t_{0}:t-1}={}_{k}\mathbf{y}^{0}_{t_{0}:t-1}\oplus\mathbf{s}_{k}\oplus M_{k}$ (9)
[…]
Since coastal inundation is expected to be correlated with tide cycles, we use hour of day, and day of month as temporal features $\mathbf{z}_{t}$. We use sinusoidal encoding (vaswani2017attention) to capture the cyclic nature of these temporal features.
${}_{k}\mathbf{g}_{t_{0}:t-1}=\textsc{sinusoid}{}({}_{k}\mathbf{z}_{t_{0}:t-1})$ (11)
@@
[…]
An overview of Diff-Sparse architecture is provided in Figure 1. We approximate (8) by the model in (13). Here, $\theta$ comprises of the trainable parameters of the convolution blocks ($\theta_{1}$), fully connected layer ($\theta_{2}$), and the denoising diffusion model ($\theta_{3}$).